# Characterization of continuous stationary fields as generalized Ornstein-Uhlenbeck fields via multi-parameter Langevin equation and multiple Riemann-Stieltjes integration

### v1

@@ 1. Introduction
Random phenomena evolving in time are often modeled using stochastic processes $X=\{X(t)\}_{t\in T}$, where $T$ is a subset of the one-dimensional Euclidean space, and the role of modeling is to assess the statistical characteristics of the process $X$. Important special classes of such processes include stationary processes, self-similar processes, and processes with stationary increments, as each of these classes allows powerful tools for modeling and understanding random phenomena across diverse fields. Indeed, stationary processes model phenomena where statistical properties remain unchanged over time, self-similar processes model phenomena where statistical characteristics are consistent across different scales (for details on self-similar processes, see the monograph [embrechts2002selfsimilar]), and stationary increment processes are particularly suitable models, for example, when considering random movements.
Connections between the three classes, stationary processes, self-similar processes, and processes with stationary increments, are already rather well-understood in the literature. For the first results relating different classes, Lamperti proved in [Lamperti-1962] that there is a one-to-one correspondence between stationary processes and self-similar processes, and the underlying bijection is nowadays known as the Lamperti transformation. Later, stationary increment processes were connected in [viitasaari2016], where (continuous) stationary processes in continuous time were characterized as solutions to the Langevin equation
@@
where $G$ is a suitable process with stationary increments, and the stochastic differential equation can be understood via integration in the Riemann-Stieltjes sense. Equation (1) gives rise to generalized Ornstein-Uhlenbeck processes, links all three classes intimately together, and one can rather freely move from one class to another with suitable transformations. Such results have later been extended to discrete time [voutilainen2017model] and to vector-valued processes both in continuous time [voutilainen2021vector] and in discrete time [voutilainen2020modeling].
@@
The features mentioned above, such as consistency of statistical properties over time or across different scales naturally appear also in the case of random fields, where the parameter $t$ of the process $\{X(t)\}_{t\in T}$ is multivariate. In this context, stationarity is still naturally defined, whereas the concepts of self-similarity and stationarity of increments are more complex. Increments of random fields are typically defined through rectangular increments, which capture changes related to all directions of the parameter space $T$ simultaneously. For literature related to stationarity of (rectangular) increments, see, e.g., [makogin2015example, makogin2019gaussian] and the references therein. A version of the Lamperti theorem for random fields was established in [genton2007self], with self-similarity defined componentwise, allowing anisotropic scaling properties required in certain applications. For more details, see e.g. [bierme2007operator], where the authors adapt an alternative approach to anisotropy via the so-called operator scaling stable random fields. For results connecting all three classes of fields via Langevin-type dynamics, we can only mention [discrete] studying fields indexed by $\mathbb{N}^{N}$ and taking values in $\mathbb{R}$, and later extensions in [voutilainen2025one] to cover fields indexed by $\mathbb{N}^{N}$ and taking values in $\mathbb{R}^{n}$. The main results of these articles provide natural dynamics, similar to the process case [voutilainen2017model], linking all three classes.
@@
In this article, we extend the results of [discrete] to fields $\{X(t)\}_{t\in\mathbb{R}^{N}}$. Compared to [discrete], the main difficulty is that differences are replaced by differentials, and the notion of Riemann-Stieltjes integrals is much more subtle in the multivariate case. In particular, for our purpose, we need to extend the notion of multiple Riemann-Stieltjes integrals over non-hyperrectangles that are non-classical domains. Some of our main contributions include several results on multiple Riemann-Stieltjes integrals, which are particularly well-suited to our framework. On top of that, we provide an approach to define multiple Riemann-Stieltjes integrals over certain non-classical domains in situations where one cannot simply apply zero-extensions of functions and use a more classical approach. With this at hand, we provide natural transformations between stationary fields, fields with stationary rectangular increments, and self-similar fields, cf. Theorem 2.7, Theorem 4.3, and Corollary 4.4. We introduce a generalization of the Langevin dynamics (1) (Definition 2.21) and, through the transformations, characterize stationary fields as solutions to these dynamics; cf. Theorem 4.5 and Theorem 4.6 (see also Theorem 4.10), providing a natural extension of the main result of [Viitasaari-2016a] from random processes to random fields. We also note that when the generalization of (1) is driven by anisotropic fractional Brownian sheets, the resulting fields can naturally be termed fractional Ornstein–Uhlenbeck fields (of the first kind).
@@
Let us also briefly discuss the existing literature on multiple Riemann-Stieltjes integration. In the univariate case, it is well-known that the Riemann-Stieltjes integral $\int f(s)dg(s)$ exists provided that one of the functions $f$ and $g$ is continuous and the other is of bounded variation. Multiple Riemann-Stieltjes integrals (over standard hyperrectangles) are well-studied in the literature, with a particular focus on the bivariate case. For details, we refer to the monographs [sard1963linear, hildebrandt1963introduction]. In the multivariate case, the assumption on bounded variation is commonly replaced by bounded variation in the sense of Hardy-Krause, originally introduced in [hardy1906double]. Different variations in the bivariate case are studied in [clarkson1933definitions]. We also cite more recent works [owen2005multidimensional, prause2015sequential] considering the general multivariate setting.
@@ 2. Preliminaries
corresponds to the increment of $f$ on the subrectangle $s_{u},t_{u}]$. Note that $\left[f(r)\right]_{s_{u}}^{t_{u}}$ is a function of N|u|N-|u| variables $r_{-u}$. We will return to this in Definition 2.8. Lastly, we present an elementary lemma on sums of binomial coefficients that we are going to utilize on several occasions. A proof can be found for example in [discrete].
@@ 2.1. Stationarity, self-similarity, and stationarity of increments
By a simple change of variables, Definitions 2.4 and 2.5 can be shown to be equivalent to the corresponding definitions in [genton2007self].
For a proof of the next theorem, we refer to [genton2007self].
@@
Furthermore, (3) is the adapted version of the definition of stationarity of increments discussed, e.g., in [makogin2019gaussian].
@@ 2.2. Multiple Riemann-Stieltjes integrals
exists as a finite number, then we call it the (unrestricted) Riemann-Stieltjes integral of $g$ with respect to $f$. We note that if the tags are of the form $\xi_{i}=(\xi^{(1)}_{i_{1}},\dots,\xi^{(N)}_{i_{N}})$, then the integral is known as the restricted Riemann-Stieltjes integral in the literature (for details in the bivariate case, see e.g. [sard1963linear]). Throughout, in our setting, we consider unrestricted Riemann-Stieltjes integrals. Moreover, we denote the integral (4) as
@@
If the rectangular domain $[s,t]$ is clear from the context, we may simply write $f\in BV$ and $f\in BVHK$. The variation in the sense of Hardy and Krause is the sum of variations in the sense of Vitali over all subrectangles $s_{v},t_{v}]$, where the coordinates v-v have been fixed to $t_{-v}$. In the one-dimensional setting, the two variations are equal to the usual variation of univariate functions. Moreover, in Definition 2.14, replacing $t_{-v}$ with any fixed point from the subrectangle $s_{-v},t_{-v}]$, would not change the class $BVHK$ of functions. See [owen2005multidimensional], [sard1963linear], and [youngfourier].
For details on the next lemma, see, e.g., [owen2005multidimensional] and the references therein.
@@
The next lemma shows that $N$ times continuously differentiable functions are bounded variations in the sense of Hardy and Krause. For a proof, we refer to [owen2005multidimensional].
@@
It was shown in [clarkson1933double] that the unrestricted integral $\int_{s}^{t}gdf$ exists for all continuous $g$ if and only if $f\in BV$ (See also [sard1963linear]). While the result was proved only in the bivariate case, it is a routine book-keeping exercise (noted in [clarkson1933double]) to obtain the claim in arbitrary dimensions. Hence our claim follows directly by the very definition of $f$ being bounded variation in the sense of Hardy and Krause. Indeed, if $f$ is bounded variation in the sense of Hardy and Krause, it implies that all lower dimensional projections (for which some variables are fixed) are of bounded variation as well. ∎
@@
For the existence of the integral (6) over the whole hyperrectangle, i.e. $v=\{1,\dots,N\}$, a sufficient condition is $f\in BV[s,t]$. A proof is presented in [prause2015sequential].
The following integration by parts formula was originally presented in [young1917multiple].
@@
Definition 2.21 is a generalization of Langevin equation to cover arbitrary $N$. For the case $N=2$, similar Langevin equations can be found at least from [de2012least] and [terdik2005notes]. In [de2012least] the parameter is a scalar and the equation does not contain mixed differentials of type $d_{t_{1}}X(t)dt_{2}$, while in [terdik2005notes] the authors consider the same equation as us except that the signs of the mixed terms are negative.
@@ 3.1. On integration over hyperrectangles
depends on all $N$ coordinates only for $w=\emptyset$. In all other cases, the square increments are identically zero (see also Theorem 2.2.5 in [prause2015sequential]) and thus, we obtain
@@ 3.2. Integration over non-hyperrectangles
$T=T(t)=\{x\in\mathbb{R}^{N}:x_{l}\leq t_{l}\text{ for all }l,\sum x_{l}\geq 0\}.$ (15)
It can be shown that
@@
$\begin{split}T\cap\mathcal{R}_{v,\emptyset}&=\{x\in\mathbb{R}^{N}:x_{v}\in[\tilde{t}_{v},t_{v}],x_{-v}=t_{-v}\}\cap\{x\in\mathbb{R}^{N}:x\leq t,\sum x_{l}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\in[\tilde{t}_{v},t_{v}],\sum x_{l}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\in[\tilde{t}_{v},t_{v}],\sum t_{-v}+\sum x_{v}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\leq t_{v},\sum t_{-v}+\sum x_{v}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{v}\in T_{v},x_{-v}=t_{-v}\},\end{split}$
@@
$T\cap\mathcal{R}_{v,w}=\{x\in\mathbb{R}^{N}:x_{w}=\tilde{t}_{w},x_{-v-w}=t_{-v-w},x_{v}\in[\tilde{t}_{v},t_{v}],\sum x_{l}\geq 0\}.$
@@
In a similar way, if $|w|\geq 2$, then $T\cap\mathcal{R}_{v,w}=\emptyset$ unless we are in the degenerate case $\sum t_{l}=0$ obtaining a set of one point. ∎
@@
as desired. Note also that when $\sum_{l}t_{l}=0$, then all the related domains are degenerate, and $\tilde{t}=t$. Since in this case $\int_{R}fdg=0$ and $\int_{T}fdg=0$ by the definitions, then also $\int_{R\setminus T}fdg=0$ by the additivity property shown above.
@@
We note also that, by the continuity of the involved functions and the fact that $\lim_{\sum_{l}t_{l}\to 0}\tilde{t}\to t$, we obtain directly from the definitions that
$\lim_{\sum_{l}t_{l}\to 0}\int_{T}f(u)dg(u)=0=\lim_{\sum_{l}t_{l}\to 0}\int_{R}f(u)dg(u)$
implying also that $\lim_{\sum_{l}t_{l}\to 0}\int_{R\setminus T}f(u)dg(u)\to 0$. Therefore, our integrals are continuous with respect to $T=T(t)$ also on the hyperplane $\sum_{l}t_{l}=0$.
@@ 4. Characterization of continuous stationary fields
The first two equalities follow from the fact that $u$ is a constant as the total differential is taken with respect to $t$ together with Theorem 3.3 (see also Theorem 2.2.5 in [prause2015sequential]). In addition, by Lemma 3.4,
@@
We end the paper by considering the uniqueness of the solution to the Langevin equation for a given initial condition. It turns out that the solution is unique only up to certain modifications.
@@
We say that processes $X(t)$ and $Y(t)$ are equivalent with respect to the parameter $\Theta$ if the process
@@
satisfies $Z]_{s}^{t}=0$ for all $s,t\in\mathbb{R}^{N}$.
@@
for some processes $H^{(u)}(t_{u})$, where now $u$ is a proper subset of $\{1,\ldots,N\}$. Note that this definition is natural, as we cannot distinguish processes $H^{(u)}(t_{u})$ in the multiple Riemann-Stieltjes sense. Indeed, their rectangular increments equal zero, since each $H^{(u)}(t_{u})$ does not depend on all the variables, and are thus essentially lower-dimensional objects. Note that, in the one-dimensional case, we would have
@@
from which the initial condition $X(0)=Y(0)$ would give uniqueness in the classical sense, i.e. $X(t)=Y(t)$ almost surely for all $t$.
We also need the following elementary lemma.
Lemma 4.8.
Suppose that $f$ solves
$\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}f_{t_{-u}}(t)=0.$
Then $f$ is given by
$f(t)=e^{-\Theta\cdot t}\sum_{\begin{subarray}{c}u\subseteq\{1,\dots,N\}\\ |u|\neq N\end{subarray}}h^{(u)}(t_{u}).$
Proof.
Define $\tilde{f}(t)=e^{\Theta\cdot t}f(t)$. We proceed by induction. Equation
$\frac{\partial}{\partial t_{1}}\tilde{f}(t)=e^{\Theta\cdot t}(\theta_{1}f(t)+f_{t_{1}}(t))$
provides the base case. We make the induction hypothesis
$\tilde{f}_{t_{\{1,\dots,j\}}}(t)=e^{\Theta\cdot t}\sum_{u\subseteq\{1,\dots,j\}}\prod_{i\in u}\theta_{i}f_{t_{\{1,\dots,j\}-u}}(t).$
Now
$\begin{split}\frac{\partial}{\partial t_{j+1}}\tilde{f}_{t_{\{1,\dots,j\}}}(t)&=e^{\Theta\cdot t}\sum_{u\subseteq\{1,\dots,j\}}\left(\prod_{i\in u}(\theta_{i})\theta_{j+1}f_{t_{\{1,\dots,j\}-u}}(t)+\prod_{i\in u}(\theta_{i})f_{t_{\{1,\dots,j+1\}-u}}(t)\right)\\ &=e^{\Theta\cdot t}\sum_{u\subseteq\{1,\dots,j+1\}}\prod_{i\in u}\theta_{i}f_{t_{\{1,\dots,j+1\}-u}}(t),\end{split}$
completing the induction step. In particular, for $N$, this leads to
$\tilde{f}_{t}(t)=e^{\Theta\cdot t}\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}f_{t_{-u}}(t)=0.$
Integrating with respect to all variables leads to
$\tilde{f}(t)=\sum_{\begin{subarray}{c}u\subseteq\{1,\dots,N\}\\ |u|\neq N\end{subarray}}h^{(u)}(t_{u}).$
This completes the proof. ∎
Lemma 4.9.
For every fixed ”initial condition” $X(0)=Z$, where $Z$ is some measurable random variable, the Langevin equation (8) admits a unique solution in the sense of Definition 4.7.
Proof.
With direct computations, we see that
$Z+e^{-\Theta\cdot t}\int_{0}^{t}e^{\Theta\cdot u}dG(u)$
is a solution, and hence it suffices to prove the uniqueness. Suppose that $X^{1}$ and $X^{2}$ are two solutions and set $X=X^{1}-X^{2}$. Then, from the linearity of the equation, we obtain that $X$ solves
$\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}d_{-u}X(t)dt_{u}=0$
with $X(0)=0$. That is, $X$ solves a partial differential equation  [^1] Note that the partial differential equation is solved in a weak sense a priori. However, since the PDE is linear with constant coefficients, it is a classical result that weak solutions are strong solutions, and hence we obtain that $X$ is in fact a (sufficiently) smooth solution.
$\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}\frac{d}{dt_{-u}}X(t)=0$
Now by Lemma 4.8 we obtain that
$X(t)=X^{1}(t)-X^{2}(t)=e^{-\Theta\cdot t}\sum_{\begin{subarray}{c}u\subseteq\{1,\dots,N\}\\ |u|\neq N\end{subarray}}H^{(u)}(t_{u}).$
@@
Let $\Theta\in(0,\infty)^{N}$. Then, a field $X=\{X(t)\}_{t\in\mathbb{R}^{N}}$ is stationary if and only if it is the unique solution, in the sense of Definition 4.7, to the Langevin equation (8) for $G\in\mathcal{G}_{\Theta,0}$ with the ”initial condition” $X(0)=\int_{-\infty}^{0}e^{\Theta\cdot u}dG(u)$. That is,

### v2

@@ 1. Introduction
Random phenomena evolving in time are often modeled using stochastic processes $X=\{X(t)\}_{t\in T}$, where $T$ is a subset of the one-dimensional Euclidean space, and the role of modeling is to assess the statistical characteristics of the process $X$. Important special classes of such processes include stationary processes, self-similar processes, and processes with stationary increments, as each of these classes allows powerful tools for modeling and understanding random phenomena across diverse fields. Indeed, stationary processes model phenomena where statistical properties remain unchanged over time, self-similar processes model phenomena where statistical characteristics are consistent across different scales (for details on self-similar processes, see the monograph [5]), and stationary increment processes are particularly suitable models, for example, when considering random movements.
Connections between the three classes, stationary processes, self-similar processes, and processes with stationary increments, are already rather well-understood in the literature. For the first results relating different classes, Lamperti proved in [9] that there is a one-to-one correspondence between stationary processes and self-similar processes, and the underlying bijection is nowadays known as the Lamperti transformation. Later, stationary increment processes were connected in [16], where (continuous) stationary processes in continuous time were characterized as solutions to the Langevin equation
@@
where $G$ is a suitable process with stationary increments, and the stochastic differential equation can be understood via integration in the Riemann-Stieltjes sense. Equation (1) gives rise to generalized Ornstein-Uhlenbeck processes, links all three classes intimately together, and one can rather freely move from one class to another with suitable transformations. Such results have later been extended to discrete time [19] and to vector-valued processes both in continuous time [18] and in discrete time [21].
@@
The features mentioned above, such as consistency of statistical properties over time or across different scales naturally appear also in the case of random fields, where the parameter $t$ of the process $\{X(t)\}_{t\in T}$ is multivariate. In this context, stationarity is still naturally defined, whereas the concepts of self-similarity and stationarity of increments are more complex. Increments of random fields are typically defined through rectangular increments, which capture changes related to all directions of the parameter space $T$ simultaneously. For literature related to stationarity of (rectangular) increments, see, e.g., [10, 11] and the references therein. A version of the Lamperti theorem for random fields was established in [6], with self-similarity defined componentwise, allowing anisotropic scaling properties required in certain applications. For more details, see e.g. [1], where the authors adapt an alternative approach to anisotropy via the so-called operator scaling stable random fields. For results connecting all three classes of fields via Langevin-type dynamics, we can only mention [20] studying fields indexed by $\mathbb{N}^{N}$ and taking values in $\mathbb{R}$, and later extensions in [17] to cover fields indexed by $\mathbb{N}^{N}$ and taking values in $\mathbb{R}^{n}$. The main results of these articles provide natural dynamics, similar to the process case [19], linking all three classes.
@@
In this article, we extend the results of [20] to fields $\{X(t)\}_{t\in\mathbb{R}^{N}}$. Compared to [20], the main difficulty is that differences are replaced by differentials, and the notion of Riemann-Stieltjes integrals is much more subtle in the multivariate case. In particular, for our purpose, we need to extend the notion of multiple Riemann-Stieltjes integrals over non-hyperrectangles that are non-classical domains. Some of our main contributions include several results on multiple Riemann-Stieltjes integrals, which are particularly well-suited to our framework. On top of that, we provide an approach to define multiple Riemann-Stieltjes integrals over certain non-classical domains in situations where one cannot simply apply zero-extensions of functions and use a more classical approach. With this at hand, we provide natural transformations between stationary fields, fields with stationary rectangular increments, and self-similar fields, cf. Theorem 2.7, Theorem 4.3, and Corollary 4.4. We introduce a generalization of the Langevin dynamics (1) (Definition 2.21) and, through the transformations, characterize stationary fields as solutions to these dynamics; cf. Theorem 4.5 and Theorem 4.6 (see also Theorem 4.10), providing a natural extension of the main result of [16] from random processes to random fields. We also note that when the generalization of (1) is driven by anisotropic fractional Brownian sheets, the resulting fields can naturally be termed fractional Ornstein–Uhlenbeck fields (of the first kind).
@@
Let us also briefly discuss the existing literature on multiple Riemann-Stieltjes integration. In the univariate case, it is well-known that the Riemann-Stieltjes integral $\int f(s)dg(s)$ exists provided that one of the functions $f$ and $g$ is continuous and the other is of bounded variation. Multiple Riemann-Stieltjes integrals (over standard hyperrectangles) are well-studied in the literature, with a particular focus on the bivariate case. For details, we refer to the monographs [14, 8]. In the multivariate case, the assumption on bounded variation is commonly replaced by bounded variation in the sense of Hardy-Krause, originally introduced in [7]. Different variations in the bivariate case are studied in [2]. We also cite more recent works [12, 13] considering the general multivariate setting.
@@ 2. Preliminaries
corresponds to the increment of $f$ on the subrectangle $s_{u},t_{u}]$. Note that $\left[f(r)\right]_{s_{u}}^{t_{u}}$ is a function of N|u|N-|u| variables $r_{-u}$. We will return to this in Definition 2.8. Lastly, we present an elementary lemma on sums of binomial coefficients that we are going to utilize on several occasions. A proof can be found for example in [20].
@@ 2.1. Stationarity, self-similarity, and stationarity of increments
By a simple change of variables, Definitions 2.4 and 2.5 can be shown to be equivalent to the corresponding definitions in [6].
For a proof of the next theorem, we refer to [6].
@@
Furthermore, (3) is the adapted version of the definition of stationarity of increments discussed, e.g., in [11].
@@ 2.2. Multiple Riemann-Stieltjes integrals
exists as a finite number, then we call it the (unrestricted) Riemann-Stieltjes integral of $g$ with respect to $f$. We note that if the tags are of the form $\xi_{i}=(\xi^{(1)}_{i_{1}},\dots,\xi^{(N)}_{i_{N}})$, then the integral is known as the restricted Riemann-Stieltjes integral in the literature (for details in the bivariate case, see e.g. [14]). Throughout, in our setting, we consider unrestricted Riemann-Stieltjes integrals. Moreover, we denote the integral (4) as
@@
If the rectangular domain $[s,t]$ is clear from the context, we may simply write $f\in BV$ and $f\in BVHK$. The variation in the sense of Hardy and Krause is the sum of variations in the sense of Vitali over all subrectangles $s_{v},t_{v}]$, where the coordinates v-v have been fixed to $t_{-v}$. In the one-dimensional setting, the two variations are equal to the usual variation of univariate functions. Moreover, in Definition 2.14, replacing $t_{-v}$ with any fixed point from the subrectangle $s_{-v},t_{-v}]$, would not change the class $BVHK$ of functions. See [12], [14], and [22].
For details on the next lemma, see, e.g., [12] and the references therein.
@@
The next lemma shows that $N$ times continuously differentiable functions are bounded variations in the sense of Hardy and Krause. For a proof, we refer to [12].
@@
It was shown in [3] that the unrestricted integral $\int_{s}^{t}gdf$ exists for all continuous $g$ if and only if $f\in BV$ (See also [14]). While the result was proved only in the bivariate case, it is a routine book-keeping exercise (noted in [3]) to obtain the claim in arbitrary dimensions. Hence our claim follows directly by the very definition of $f$ being bounded variation in the sense of Hardy and Krause. Indeed, if $f$ is bounded variation in the sense of Hardy and Krause, it implies that all lower dimensional projections (for which some variables are fixed) are of bounded variation as well. ∎
@@
For the existence of the integral (6) over the whole hyperrectangle, i.e. $v=\{1,\dots,N\}$, a sufficient condition is $f\in BV[s,t]$. A proof is presented in [13].
The following integration by parts formula was originally presented in [23].
@@
Definition 2.21 is a generalization of the Langevin equation to cover arbitrary $N$. For the case $N=2$, similar Langevin equations can be found at least from [4] and [15]. In [4] the parameter is a scalar and the equation does not contain mixed differentials of type $d_{t_{1}}X(t)dt_{2}$, while in [15] the authors consider the same equation as us except that the signs of the mixed terms are negative.
@@ 3.1. On integration over hyperrectangles
depends on all NN coordinates only for $w=\emptyset$. In all other cases, the square increments are identically zero (see also Theorem 2.2.5 in [13]) and thus, we obtain
@@ 3.2. Integration over non-hyperrectangles
$T=T(t)=\{x\in\mathbb{R}^{N}:x_{l}\leq t_{l}\text{ for all }l,\sum x\geq 0\},$ (15)
where $\sum x$ is the sum of all elements of the vector $x$. It can be shown that
@@
$\begin{split}T\cap\mathcal{R}_{v,\emptyset}&=\{x\in\mathbb{R}^{N}:x_{v}\in[\tilde{t}_{v},t_{v}],x_{-v}=t_{-v}\}\cap\{x\in\mathbb{R}^{N}:x\leq t,\sum x\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\in[\tilde{t}_{v},t_{v}],\sum x\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\in[\tilde{t}_{v},t_{v}],\sum t_{-v}+\sum x_{v}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{-v}=t_{-v},x_{v}\leq t_{v},\sum t_{-v}+\sum x_{v}\geq 0\}\\ &=\{x\in\mathbb{R}^{N}:x_{v}\in T_{v},x_{-v}=t_{-v}\},\end{split}$
@@
$T\cap\mathcal{R}_{v,w}=\{x\in\mathbb{R}^{N}:x_{w}=\tilde{t}_{w},x_{-v-w}=t_{-v-w},x_{v}\in[\tilde{t}_{v},t_{v}],\sum x\geq 0\}.$
@@
In a similar way, if $|w|\geq 2$, then $T\cap\mathcal{R}_{v,w}=\emptyset$ unless we are in the degenerate case $\sum t=0$ obtaining a set of one point. ∎
@@
as desired. Note also that when $\sum t=0$, then all the related domains are degenerate, and $\tilde{t}=t$. Since in this case $\int_{R}fdg=0$ and $\int_{T}fdg=0$ by the definitions, then also $\int_{R\setminus T}fdg=0$ by the additivity property shown above.
@@
We also note that, by the continuity of the involved functions and the fact that $\lim_{\sum t\to 0}\tilde{t}\to t$, we obtain directly from the definitions that
$\lim_{\sum t\to 0}\int_{T}f(u)dg(u)=0=\lim_{\sum t\to 0}\int_{R}f(u)dg(u)$
implying also that $\lim_{\sum t\to 0}\int_{R\setminus T}f(u)dg(u)\to 0$. Therefore, our integrals are continuous with respect to $T=T(t)$ also on the hyperplane $\sum t=0$.
@@ 4. Characterization of continuous stationary fields
The first two equalities follow from the fact that $u$ is a constant as the total differential is taken with respect to $t$ together with Theorem 3.3 (see also Theorem 2.2.5 in [13]). In addition, by Lemma 3.4,
@@
We end the paper by considering the uniqueness of the solution to the Langevin equation. It turns out that the solution is unique up to certain lower-dimensional components that can be eliminated by posing appropriate ”boundary conditions”.
@@
We say that fields $X(t)$ and $Y(t)$ are equivalent with respect to the parameter $\Theta$ if the field
@@
satisfies $Z]_{s}^{t}=0$ almost surely for all $s,t\in\mathbb{R}^{N}$.
@@
[…]
Lemma 4.8.
The Langevin equation (8) admits a unique solution in the sense of Definition 4.7.
Proof.
Suppose that $X$ and $\tilde{X}$ are two solutions and set $Z=X-\tilde{X}$. Then, from the linearity of the equation, we obtain that $Z$ solves
$\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}d_{-u}Z(t)dt_{u}=0.$
Multiplying with $e^{\Theta\cdot t}$ yields
$\begin{split}\sum_{u\subseteq\{1,\dots,N\}}\prod_{i\in u}\theta_{i}e^{\Theta\cdot t}d_{-u}Z(t)dt_{u}&=\sum_{u\subseteq\{1,\dots,N\}}\frac{de^{\Theta\cdot t}}{dt_{u}}d_{-u}Z(t)dt_{u}=d(e^{\Theta\cdot t}Z(t))=0\end{split}$
by Theorem 3.3 under the standing assumption of almost surely continuous realizations. This gives $e^{\Theta\cdot r}Z(r)]_{s}^{t}=0$ for all $s,t$, and as in the proof of Theorem 4.3,
$e^{\Theta\cdot t}Z(t)=\sum_{\begin{subarray}{c}u\subseteq\{1,\dots,N\}\\ |u|\neq N\end{subarray}}H^{(u)}(t_{u}).$
Remark 4.9.
By fixing some (random) ”boundary conditions”, we may eliminate some of the fields $H^{(u)}(t_{u})$. For example, if in the two-dimensional setting $X(0,t_{2})=X_{2}(t_{2})$, we obtain that
$H^{(2)}(t_{2})=-H^{(1)}(0)-C.$
From which,
[…]
Note that, in the one-dimensional case, we would have
@@
from which the condition $X(0)=Y(0)$ would give uniqueness in the classical sense, i.e. $X(t)=Y(t)$ almost surely for all $t$.
@@
Let $\Theta\in(0,\infty)^{N}$. A field $X=\{X(t)\}_{t\in\mathbb{R}^{N}}$ is stationary if and only if it is the unique solution to the Langevin equation (8) for $G\in\mathcal{G}_{\Theta,0}$ with the following ”boundary condition”: $X(t)=e^{-\Theta\cdot t}\int_{-\infty}^{t}e^{\Theta\cdot u}dG(u)$ for all $t$ such that $t_{l}=0$ at least for one $l\in\{1,\dots,N\}$. That is,
$X(t)=\left((\mathcal{L}_{\Theta}^{-1}\circ\mathcal{M}_{\Theta}^{-1})(G)\right)(t)=e^{-\Theta\cdot t}\int_{-\infty}^{t}e^{\Theta\cdot u}dG(u)\quad\text{for all }t\in\mathbb{R}^{N}.$